OpenStudy (anonymous):

Arthur is trying to find the equation of a perpendicular to y = '1 over 4 x + 4 in slope-intercept form that passes through the point (−2, 6). Which of the following equations will he use? A. y − 6 = 1 over 4 (x − (−2)) B. y − (−2) = 1 over 4 (x − 6) C. y − 6 = −4(x − (−2)) D. y − (−2) = −4(x − 6)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the given slope is 1/4 what's the perpendicular slope?

OpenStudy (anonymous):

I'm not sure.

jimthompson5910 (jim_thompson5910):

flip the fraction: 1/4 ----> 4/1 = 4 flip the sign: 4 ----> -4

jimthompson5910 (jim_thompson5910):

original slope: +1/4 perpendicular slope: -4/1 or -4

jimthompson5910 (jim_thompson5910):

does that seem familiar from the lesson?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

so m = -4

jimthompson5910 (jim_thompson5910):

the given point is (x1,y1) = (-2,6)

jimthompson5910 (jim_thompson5910):

so we plug these values into the point slope equation \[\Large y - y_1 = m(x-x_1)\] \[\Large y - 6 = -4(x-(-2))\]

OpenStudy (anonymous):

So is it D?

jimthompson5910 (jim_thompson5910):

we could keep going but we can stop here because there is an answer choice that has this

OpenStudy (anonymous):

I can't tell if it's B or D.

jimthompson5910 (jim_thompson5910):

look closely and compare what I have with the choices

jimthompson5910 (jim_thompson5910):

hopefully you see how I got that

OpenStudy (anonymous):

It seems like it's D, am I right?

jimthompson5910 (jim_thompson5910):

you mixed up 2 terms. Look closer

OpenStudy (anonymous):

So it's B?

jimthompson5910 (jim_thompson5910):

it sounds like you're randomly guessing

jimthompson5910 (jim_thompson5910):

like I said, look for the one that matches with \[\Large y - 6 = -4(x-(-2))\]

OpenStudy (anonymous):

Ohhh C....I'm sorry my internet wasn't working.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you. :)

jimthompson5910 (jim_thompson5910):

you're welcome